OpenStudy (anonymous):

differentiate logx3^x

OpenStudy (anonymous):

is this \[\log_x(3^x)\]??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i hope not because how in the world are you going to differentiate \[\log_x\]?

myininaya (myininaya):

\[\log_x(3^x)=\frac{\ln(3^x)}{\ln(x)}=\frac{x \ln(3)}{\ln(x)}\]

OpenStudy (anonymous):

oh maybe we can try this \[\log_x(3^x)=x\log_x(3)=\frac{x\ln(3)}{\ln(x)}\]

myininaya (myininaya):

lol

OpenStudy (anonymous):

what myininaya said!

myininaya (myininaya):

\[\ln(3) \cdot \frac{d}{dx} (\frac{x}{\ln(x)})\]

myininaya (myininaya):

use quotient rule

OpenStudy (anonymous):

zactly

myininaya (myininaya):

\[\ln(3) \cdot \frac{(x)' \ln(x)-x \cdot (\ln(x))'}{(\ln(x))^2}\]

OpenStudy (anonymous):

get \[\frac{\ln(x)-1}{\ln^2(x)}\]

myininaya (myininaya):

don't forget you constant multiply ln(3)

OpenStudy (anonymous):

multiply the result by \[\ln(3)\]

OpenStudy (anonymous):

i was getting there...

myininaya (myininaya):

no! :)

myininaya (myininaya):

goodnight guys

OpenStudy (anonymous):

ok you are faster g'night!